# Consistency tests between SDSS and DESI BAO measurements

### v2

@@ Abstract
In this work, we investigate whether the baryon acoustic oscillation (BAO) measurements from redshift surveys, like the Sloan Digital Sky Survey (SDSS), and the Dark Energy Spectroscopic Instrument (DESI), are consistent with each other. We do so by obtaining the Hubble and deceleration parameter, respectively $H(z)$ and $q(z)$, from both datasets using a non-parametric reconstruction, so that our results do not depend on any a priori assumptions about the underlying cosmological model. We find that the reconstructed $H(z)$ and $q(z)$ from SDSS are significantly inconsistent with those obtained from DESI, and that both are only marginally consistent with the $\Lambda$CDM model ($\sim 3\sigma$ confidence level). Interestingly, the combined SDSS and DESI dataset reconciles with the standard model. These results are mostly unchanged with respect to different assumptions on the sound horizon scale value, as well as different reconstruction kernels. We also verify the results for the null diagnostic $\mathcal{O}_{\rm m}(z)$, finding that SDSS favours a quintessence-like dark energy model, whereas a phantom-like dark energy is preferred by DESI data, and once again the combined dataset strongly agrees with $\Lambda$CDM. Therefore, our results call the attention for further examination of such inconsistency, as they can lead to biased and divergent results regarding the validity of the standard model, or the suggestion of new physics.
@@ I Introduction
In this work, we focus on a comparison between the state-of-the-art BAO measurements from DESI and the previous releases of Sloan Digital Sky Survey (SDSS). Recently, a comparison between DESI and non-DESI data has been carried out Park et al. (2024), with the latter including CMB anisotropy data and non-CMB data, where it was found that such a data compilation is better constrained to favour a flat $w_{0}w_{a}$ model, compared to a flat $\Lambda$CDM model. Our goal, however, is to verify in a model-independent way whether there is any inconsistency between the DESI and SDSS BAO datasets, as it would indicate potential systematics that might lead to discrepant conclusions about the validity of the SCM, as well as the evidence of new physics. We perform this comparison by means of non-parametric reconstructions of observable quantities, such as the cosmic expansion rate through the Hubble parameter, $H(z)$, and the deceleration parameter, $q(z)$, that can be obtained from each individual dataset and their combination. These reconstructions allow us to circumvent a priori conclusions due to the assumption of a cosmological model to describe those quantities of interest. Other model-agnostic studies of dark energy in the light of DESI have been performed in Mukhopadhayay et al. (2024); Rebouças et al. (2024); Ruchika (2024); Yang et al. (2024), and some interesting works on obtaining model-independent distances from BAO can be found in Anselmi et al. (2018); O’Dwyer et al. (2020); Anselmi et al. (2023). Interestingly, we find that the DESI and SDSS BAO datasets are significantly inconsistent with each other, and only marginally consistent with the SCM predictions, although the combined SDSS and DESI dataset is in great agreement with $\Lambda$CDM. This result is further illustrated by the null diagnostic $\mathcal{O}_{\rm m}$, which shows discrepant deviations from the SCM scenario between each individual dataset, yet again the agreement with the SCM is restored when the combined dataset is considered.
@@ III Results
The results of the null diagnostic $\mathcal{O}_{\rm m}(z)$ for both Matern(7/2) kernel and low-$z$ sound horizon values are presented in Figs. 7 and 8, respectively. In the latter case, the thick blue line corresponds to $\Omega_{\rm m}=0.334\pm 0.018$, and we assumed the prior $H_{0}=73.3\pm 1.04\;\mathrm{km}\;\mathrm{s}^{-1}\;\mathrm{Mpc}^{-1}$ to compute $\mathcal{O}_{\rm m}(z)$ and its uncertainty, i.e., the best-fitted $\Omega_{\rm m}$ and $H_{0}$ parameters, respectively, assuming a flat $\Lambda$CDM model to the PantheonPlus and SH0ES data Brout et al. (2022). While the alternative GP kernel yields similar results to the default kernel choice, the results are quite different for some specific datasets when we assume the low-$z$ sound horizon measurement and the Hubble constant. In this case, we can still see that quintessence-like and phantom-like dark energy models are favoured by SDSS and DESI datasets alone, respectively, albeit with a slightly lower significance. This is also supported by recent works Wolf et al. (2024a, b), whereby it is shown that there are no conclusive evidences to consider thawing quintessence models as the description for evolving dark energy. However, we find that the agreement between DESI+SDSS C1 data and SCM marginally worsens when the low-$z$ $r_{\rm d}$ and $H_{0}$ values are assumed (at slightly over $1\sigma$ confidence level), suggesting a quintessence-like behaviour for dark energy by the same fashion of SDSS data alone, while the opposite trend occurs for the C2 case. But even in this specific case, we can still see the same general trend of our main results, that is, the combination of DESI and SDSS data is in better agreement with the SCM predictions than the individual datasets. So, overall, the results obtained from different assumptions on the GP kernel and sound horizon scale present the same trend as in the default case. This demonstrates the robustness of our results with regards to them. Interestingly, however, recent results Jiang et al. (2024b) suggest that the upper limits on the neutrino mass sum as obtained from DESI are tightened further if quintessence or non-phantom dark energy models are considered, warranting further investigation into the preference of one class of model over another.
@@ IV Conclusion
Furthermore, the $\mathcal{O}_{\rm m}(z)$ diagnostic seems to suggest that DESI alone prefers a phantom-like dark energy model, whereas SDSS alone prefers a quintessence-like dark energy model. We further carry out our analysis with a different GP kernel, Matern72, instead of the squared exponential one, as well as impose a different sound horizon scale assuming a low-$z$ prior. Barring some minor changes, the overall behaviour of the reconstructions for all the above cases remains the same. We believe that our analysis calls for further investigation of existing results in order to address this inconsistency between SDSS and DESI. We can expect to have a better insight once the low redshift data of DESI become available and can be compared with SDSS data at similar redshifts. In the meantime, it would be worthwhile to explore these disagreements further so that biased conclusions on the validity of the standard model can be avoided.

### v3

@@ Abstract
In this work, we investigate whether the baryon acoustic oscillation (BAO) measurements from redshift surveys, like the Sloan Digital Sky Survey (SDSS), and the Dark Energy Spectroscopic Instrument (DESI), are consistent with each other. We do so by obtaining the Hubble and deceleration parameter, respectively $H(z)$ and $q(z)$, from both datasets using a non-parametric reconstruction, so that our results do not depend on any a priori assumptions about the underlying cosmological model. We find that the reconstructed $H(z)$ and $q(z)$ from SDSS are significantly inconsistent with those obtained from DESI, and that both are only marginally consistent with the $\Lambda$CDM model ($\sim 3\sigma$ confidence level). Interestingly, the combined SDSS and DESI dataset reconciles with the standard model. These results are mostly unchanged with respect to different assumptions on the sound horizon scale value, as well as different reconstruction kernels. We also verify the results for the null diagnostic $\mathcal{O}_{\mathrm{m}}(z)$, finding that they suggest different trends for dark energy models using DESI and SDSS BAO measurements, and once again the combined dataset strongly agrees with $\Lambda$CDM. Therefore, our results call the attention for further examination of such inconsistency, as they can lead to biased and divergent results regarding the validity of the standard model, or the suggestion of new physics.
@@ I Introduction
In this work, we focus on a comparison between the state-of-the-art BAO measurements from DESI and the previous releases of Sloan Digital Sky Survey (SDSS). Recently, a comparison between DESI and non-DESI data has been carried out Park et al. (2024), with the latter including CMB anisotropy data and non-CMB data, where it was found that such a data compilation is better constrained to favour a flat $w_{0}w_{a}$ model, compared to a flat $\Lambda$CDM model. Our goal, however, is to verify in a model-independent way whether there is any inconsistency between the DESI and SDSS BAO datasets, as it would indicate potential systematics that might lead to discrepant conclusions about the validity of the SCM, as well as the evidence of new physics. We perform this comparison by means of non-parametric reconstructions of observable quantities, such as the cosmic expansion rate through the Hubble parameter, $H(z)$, and the deceleration parameter, $q(z)$, that can be obtained from each individual dataset and their combination. These reconstructions allow us to circumvent a priori conclusions due to the assumption of a cosmological model to describe those quantities of interest. Other model-agnostic studies of dark energy in the light of DESI have been performed in Mukhopadhayay et al. (2024); Rebouças et al. (2024); Ruchika (2024); Yang et al. (2024), and some interesting works on obtaining model-independent distances from BAO can be found in Anselmi et al. (2018); O’Dwyer et al. (2020); Anselmi et al. (2023). Interestingly, we find that the DESI and SDSS BAO datasets are significantly inconsistent with each other, and only marginally consistent with the SCM predictions, although the combined SDSS and DESI dataset is in great agreement with $\Lambda$CDM. This result is further illustrated by the null diagnostic $\mathcal{O}_{\mathrm{m}}$, which shows discrepant deviations from the SCM scenario between each individual dataset, yet again the agreement with the SCM is restored when the combined dataset is considered.
@@ III Results
The results of the null diagnostic $\mathcal{O}_{\rm m}(z)$ for both Matern(7/2) kernel and low-$z$ sound horizon values are presented in Figs. 7 and 8, respectively. In the latter case, the thick blue line corresponds to $\Omega_{\rm m}=0.334\pm 0.018$, and we assumed the prior $H_{0}=73.3\pm 1.04\;\mathrm{km}\;\mathrm{s}^{-1}\;\mathrm{Mpc}^{-1}$ to compute $\mathcal{O}_{\rm m}(z)$ and its uncertainty, i.e., the best-fitted $\Omega_{\rm m}$ and $H_{0}$ parameters, respectively, assuming a flat $\Lambda$CDM model to the PantheonPlus and SH0ES data Brout et al. (2022). While the alternative GP kernel yields similar results to the default kernel choice, the results are quite different for some specific datasets when we assume the low-$z$ sound horizon measurement and the Hubble constant. In this case, we can still see the same trends for SDSS and DESI datasets alone, albeit with a slightly lower significance. This is also supported by recent works Wolf et al. (2024a, b), whereby it is shown that there are no conclusive evidences to consider thawing quintessence models as the description for evolving dark energy. However, we find that the agreement between DESI+SDSS C1 data and SCM marginally worsens when the low-$z$ $r_{\rm d}$ and $H_{0}$ values are assumed (at slightly over $1\sigma$ confidence level). But even in this specific case, we can still see the same general trend of our main results, that is, the combination of DESI and SDSS data is in better agreement with the SCM predictions than the individual datasets. So, overall, the results obtained from different assumptions on the GP kernel and sound horizon scale present the same trend as in the default case. This demonstrates the robustness of our results with regards to them. Interestingly, however, recent results Jiang et al. (2024b) suggest that the upper limits on the neutrino mass sum as obtained from DESI are tightened further if quintessence or non-phantom dark energy models are considered, warranting further investigation into the preference of one class of model over another.
@@ IV Conclusion
Furthermore, the $\mathcal{O}_{\rm m}(z)$ diagnostic seems to suggest that DESI and SDSS alone prefer different behaviour for dark energy between themselves. We further carry out our analysis with a different GP kernel, Matern72, instead of the squared exponential one, as well as impose a different sound horizon scale assuming a low-$z$ prior. Barring some minor changes, the overall behaviour of the reconstructions for all the above cases remains the same. We believe that our analysis calls for further investigation of existing results in order to address this inconsistency between SDSS and DESI. We can expect to have a better insight once the low redshift data of DESI become available and can be compared with SDSS data at similar redshifts. In the meantime, it would be worthwhile to explore these disagreements further so that biased conclusions on the validity of the standard model can be avoided.